A particular hybrid car travels approximately 123 mi on 3 gal of gas. Find the amount of gas required for A 574-mi trip.
The car needs ____ gallons of gas for a 574-mi trip.
(Type an interger or a decimal)

Answer by stanbon(57337)

A particular hybrid car travels approximately 123 mi on 3 gal of gas. Find the amount of gas required for A 574-mi trip.
-------
Use a proportion:
gallons/distance = gallons/distance
------------------------------
x/574 = 3/123
x = (574)(3/123) = 14 gallons
================
Cheers,
Stan H.
---------------------